Assuming a release of 173 MeV per fission reaction, calculate how many reactions must occur per second to produce a power output of 160 MW. Express your answer using two significant figures
Answer #1

Here,

energy released per fission = 173 MeV

energy released per fission = 173 * 1.602 *10^-19 * 1 *10^6

energy released per fission = 2.77 *10^-11 J

Number of reaction per sec = 160 *10^6/(2.77 *10^-11)

Number of reaction per sec = 5.77 *10^18 1/s

the Number of reaction per sec is 5.77 *1018 1/s